# Generalizing the matching decoder for the Chamon code

### v1

@@ Abstract
Different choices of quantum error-correcting codes can reduce the demands on the physical hardware needed to build a quantum computer. To achieve the full potential of a code, we must develop practical decoding algorithms that can correct errors that have occurred with high likelihood. Matching decoders are very good at correcting local errors while also demonstrating fast run times that can keep pace with physical quantum devices. We implement variations of a matching decoder for a three-dimensional, non-CSS, low-density parity check code known as the Chamon code, which has a non-trivial structure that does not lend itself readily to this type of decoding. The non-trivial structure of the syndrome of this code means that we can supplement the decoder with additional steps to improve the threshold error rate, below which the logical failure rate decreases with increasing code distance. We find that a generalized matching decoder that is augmented by a belief-propagation step prior to matching gives a threshold of $10.5\%$ for depolarising noise.
I Introduction
@@
II The Chamon code
II.1 The stabilizer group
@@
Figure 1: The Chamon code. (a) The Chamon code lattice. Every white vertex represents a qubit and each black vertex represents a stabilizer. (b) A stabilizer of the Chamon code. All stabilizers in the Chamon code are identical and apply a Pauli-X operator on the two neighboring qubits in the x axis, a Pauli-Y operator on the two neighboring qubits in the y axis, and similarly a Pauli-Z for the z axis. (c) Examples of single-qubit Pauli errors, together with their violated stabilizer operators.
II.2 The symmetries of the Chamon code
@@
where the plane lies perpendicular to the vector $\mathbf{r}=(r_{x},r_{y},r_{z})$ with $r_{x},r_{y},r_{z}=\pm 1$ and $C$ is a constant that translates the plane along $\mathbf{r}$. Given that $-\mathbf{r}$ is parallel to $\mathbf{r}$, we have four unique orientations of plane by this definition. One can check that for each of these subsets, the product of all the elements of this subset gives identity. It follows from this observation that any Pauli error will always produce an even number of defects among the stabilizers of a code symmetry. In fact, one can check that errors always produce defects in pairs among each of these subsets.
@@
III A matching decoder for the Chamon code
Error correction requires a decoder to attempt to correct the error that has occurred. The decoder uses the syndrome to identify the recovery operation which brings the state back to its original logical state with high probability. We will assume the noise is a Pauli noise for simplicity, as the stabilizer measurements project a general noise channel to a noise close to Pauli noise [46, 47, 48]. For some specific instance of a Pauli error $E$, which specifies which qubits suffered from which single qubit Pauli noise, the stabilizer measurement results in a syndrome $s$. The decoder takes the syndrome $s$ as input and returns as output a recovery operation $C$ which it deems most probable to bring the state back to its encoded logical state. When analyzing the logical error rate of a code via simulations where $E$ is known, we can determine whether or not the recovery operation indeed succeeded in bringing the state back to the original logical state by observing whether $CE$ is a logical operator, in which case the decoding failed and resulted in a logical failure, or whether $CE$ is in the stabilizer group, in which case the decoding succeeded.
Here we describe the decoding algorithms we implement for the Chamon code. All the decoders we implement are based on a matching step on symmetries, that is followed by a local correction once matching is completed. We describe this first. We then go on to describe modifications to the basic implementation of the matching algorithm. We consider a greedy initialisation step, as well as a belief propagation step to improve the accuracy of the MWPM subroutine.
III.1 Basic matching
The decoder uses matching subroutines to identify small clusters of syndrome defects that can be corrected locally. The clusters are obtained by connecting defects that share an edge returned from matching subroutines, where matching is performed on all planar subsets of stabilizers as defined in Eqn. (2). A pair of defects are included in a common cluster if they share an edge on any matching subroutine. In general, a cluster contains many defects, as each defect will share edges with multiple defects over all of the matching subroutines.
@@ III.1 Basic matching
III.1.1 Step 1 - Matching on symmetries
On each of the symmetries described in Sec. II.2, every single-qubit Pauli error gives rise to up to two syndrome defects, and so can be decoded using MWPM, see Fig. 3. The first step of the matching decoder then is to apply matching, implemented using PyMatching [13], on each of the code symmetries, as in Eqn. (2), separately. Since we are using these decoders as a first step, in a multi-step decoder, we are not interested in the specific qubits it deems should be flipped. Instead, we take as output the previous step in the decoder – the pairs of syndrome defects it deems should be connected. And so the output of this first step in the decoding process is a set of connections between the syndrome defects.
@@ III.1.1 Step 1 - Matching on symmetries
III.1.2 Step 2 - dividing syndrome defects into clusters
@@
III.1.3 Step 3 - correcting each cluster via sweeping decoder
@@
III.2 Greedy matching
We find an improvement on the basic version of the decoder by adding an initial greedy step to the decoding algorithm. In another context, this greedy step could be interpreted as initialization [38] or predecoding [39], or an initial stage using a small-set-flip decoder [50].
@@ III.3 Belief matching
IV Simulation and results
The threshold of a QECC is one method to determine how well the code overcomes noise. It describes a family of codes characterized by a parameter $L$. The threshold is the noise rate $p_{*}$ under which the logical error rate can be suppressed arbitrarily by taking a representative code from the family with sufficiently large length $L$ [53]. This is not to be confused with another definition of threshold, sometimes also called a pseudo-threshold, which relates the logical error of a code to the physical noise [1].
(a)Refer to caption (b)Refer to caption
@@ IV Simulation and results
Let us finally compare these decoders directly by concentrating on the logical error rates of a code of fixed distance, see Fig.  7. Here we show the performance of the decoders we have proposed together with a standard implementation of the belief-propagation ordered statistics decoder (BP-OSD) [54, 51, 55]. This decoder is well studied in the literature due to the fact that it is readily applicable to any code where the check matrix is known. We find that below threshold the belief matching on symmetries decoder performs best of all decoders. We will also note that its runtime was significantly faster than BP-OSD. Given that both decoders begin using a belief-propagation step, our results indicate that matching subroutines may provide a better option than ordered statistics decoding when interpreting the results of belief propagation for cases where the message passing step does not converge.
V Summary
@@
Acknowledgements.

### v2

@@ Abstract
Different choices of quantum error-correcting codes can reduce the demands on the physical hardware needed to build a quantum computer. To achieve the full potential of a code, we must develop practical decoding algorithms that can correct errors that have occurred with high likelihood. Matching decoders are very good at correcting local errors while also demonstrating fast run times that can keep pace with physical quantum devices. We implement variations of a matching decoder for a three-dimensional, non-CSS, low-density parity check code known as the Chamon code, which has a non-trivial structure that does not lend itself readily to this type of decoding. The non-trivial structure of the syndrome of this code means that we can supplement the decoder with additional steps to improve the threshold error rate, below which the logical failure rate decreases with increasing code distance. We find that a generalized matching decoder that is augmented by a belief-propagation step prior to matching gives a threshold of $10.5\%$ for depolarizing noise.
1 Introduction
@@
2 The Chamon code
2.1 The stabilizer group
@@
Figure 1: The Chamon code. (a) The Chamon code lattice. Every white vertex represents a qubit and every black vertex represents a stabilizer. (b) A stabilizer of the Chamon code. All stabilizers in the Chamon code are identical and apply a Pauli-X operator on the two neighboring qubits in the x axis, a Pauli-Y operator on the two neighboring qubits in the y axis, and similarly a Pauli-Z for the z axis. (c) Examples of single-qubit Pauli errors, together with their violated stabilizer operators.
2.2 The symmetries of the Chamon code
@@
where the plane lies perpendicular to the vector $\mathbf{r}=(r_{x},r_{y},r_{z})$ with $r_{x},r_{y},r_{z}=\pm 1$ and $C$ is a constant that translates the plane along $\mathbf{r}$. Given that $-\mathbf{r}$ is parallel to $\mathbf{r}$, we have four unique plane orientations by this definition. There are $d/2$ different symmetries in each orientation, and so a total of $2d$ symmetries. One can check that for each of these subsets, the product of all the elements of this subset gives identity. It follows from this observation that any Pauli error will always produce an even number of defects among the stabilizers of a code symmetry. In fact, one can check that errors always produce defects in pairs among each of these subsets.
@@
3 A matching decoder for the Chamon code
Error correction requires a decoder to attempt to correct the error that has occurred. The decoder uses the syndrome to identify the recovery operation that brings the state back to its original logical state with high probability. We will assume the noise is Pauli noise for simplicity, as the stabilizer measurements project a general noise channel to a noise close to Pauli noise [46, 47, 48]. For some specific instance of a Pauli error $E$, which specifies which qubits suffered from which single qubit Pauli noise, the stabilizer measurement results in a syndrome $s$. The decoder takes the syndrome $s$ as input and returns as output a recovery operation $C$ which it deems most probable to bring the state back to its encoded logical state. When analyzing the logical error rate of a code via simulations where $E$ is known, we can determine whether or not the recovery operation indeed succeeded in bringing the state back to the original logical state by observing whether $CE$ is a logical operator, in which case the decoding failed and resulted in a logical failure, or whether $CE$ is in the stabilizer group, in which case the decoding succeeded.
Here we describe the decoding algorithms that we implement for the Chamon code, with the source code available at [49]. All the decoders we implement are based on a matching step on symmetries that is followed by a local correction once matching is completed. We describe this first. We then go on to describe modifications to the basic implementation of the matching algorithm. We consider a greedy initialization step, as well as a belief propagation initialization step to improve the accuracy of the MWPM subroutine.
3.1 Basic matching
The decoder uses matching subroutines to identify small clusters of syndrome defects that can be corrected locally. The clusters are obtained by connecting defects that share an edge returned from the matching subroutines, where the matching is performed on all planar subsets of stabilizers as defined in Eqn. (2). A pair of defects are included in a common cluster if they share an edge on any matching subroutine. In general, a cluster contains many defects, as each defect will share edges with multiple defects over all of the matching subroutines.
@@ 3.1 Basic matching
3.1.1 Step 1 - Matching on symmetries
On each of the symmetries described in Sec. 2.2, every single-qubit Pauli error gives rise to up to two syndrome defects, and so can be decoded using MWPM, see Fig. 3. The first step of the matching decoder then is to apply matching, implemented using PyMatching [13], on each of the code symmetries, as in Eqn. (2), separately. Since we are using these decoders as a first step, in a multi-step decoder, we are not interested in the specific qubits it deems should be flipped. Instead, we take as output the previous step in the decoder – the pairs of syndrome defects it deems should be connected. And so the output of this first step in the decoding process is a set of connections between the syndrome defects.
@@ 3.1.1 Step 1 - Matching on symmetries
3.1.2 Step 2 - dividing syndrome defects into clusters
@@
3.1.3 Step 3 - correcting each cluster via sweeping decoder
@@
3.2 Greedy matching
We suggest an improvement on the basic version of the decoder by adding an initial greedy step to the decoding algorithm. In another context, this greedy step could be interpreted as initialization [38] or predecoding [39], or an initial stage using a small-set-flip decoder [51].
@@ 3.3 Belief matching
4 Simulation and results
The threshold of a QECC is one method to determine how well the code overcomes noise. It is a property of a family of codes characterized by a parameter $L$. The threshold is the noise rate $p_{*}$ under which the logical error rate can be suppressed arbitrarily by taking a representative code from the family with sufficiently large length $L$ [54]. This is not to be confused with another definition of threshold, sometimes also called a pseudo-threshold, which relates the logical error of a code to the physical noise [1].
@@ 4 Simulation and results
Let us finally compare these decoders directly by concentrating on the logical error rates of a code of fixed distance, see Fig.  7. Here we show the performance of the decoders we have proposed together with a standard implementation of the belief-propagation ordered statistics decoder (BP-OSD) [55, 52, 56]. This decoder is well studied in the literature due to the fact that it is readily applicable to any code where the check matrix is known. We find that below threshold the belief matching on symmetries decoder performs best of all decoders. We will also note that its runtime was significantly faster than BP-OSD. The complexity of the BP-OSD decoder is $O(n^{3})$ [55]. On the other hand, the worst-scaling part of the belief matching decoder is the MWPM subroutine which scales as $\tilde{O}(n^{3})$ [13]. We run this subroutine on symmetries which include $O(n^{2/3})$ stabilizers and have $O(n^{1/3})$ such symmetries, and so in total the complexity of the belief matching decoder is $\tilde{O}(n^{7/3})$. Given that both decoders begin using a belief-propagation step, our results indicate that matching subroutines may provide a better option than ordered statistics decoding when interpreting the results of belief propagation for cases where the message passing step does not converge.
5 Summary
@@
A central challenge in this work, and the source of its most compelling open question, lies in the difficulty of identifying suitable symmetry planes on which the matching procedure can be applied. These symmetry planes must meet two key criteria: (1) the matching decoder must be implementable on each symmetry plane, and (2) the clusters produced by the decoding processes on the symmetry planes must be independently correctable. Because of these constraints, identifying appropriate symmetry planes is a highly non-trivial task and, in this work, was carried out manually. This raises an important open question: is it possible to develop a general algorithm or systematic method for identifying such symmetry planes for arbitrary codes?
Acknowledgements